OpenStudy (anonymous):

Eric is making a scale drawing of the ground floor of his house. If the scale is 2 in = 5 ft, what number of inches will he use for a length of 55 ft? Is It 22 in. ?

OpenStudy (anonymous):

yes, set up a ratio 2/x = 5/55 cross multiply 2(55) = 5(x) 110 = 5(x) divide both sides by 5 22 = x as you can see, 22in. is correct

OpenStudy (anonymous):

thank you [:

OpenStudy (anonymous):

ur welcome